The angular speed of an automobile engine is increased at a constant rate from 1470 rev/min to 3320 rev/min in 8.16 s. (a) What is its angular acceleration in revolutions per minute-squared? (b) How many revolutions does the engine make during this 8.16 s interval?
Answer #1

(a)the angular acceleration is calculated as follows:

ωf = ωi + αt
      α = (ωfi)/t = (3320 - 1470)/(8.16/60) = 13602.94 rev/min2 = 13.6x10^3 rev/min^2

(b)

Number of revolutions made by the engine in8.16 s is,

θ =(1/2)( ωif)t =(1/2)(1470+3320)(8.16/60) = 325.72 rev